OpenStudy (anonymous):

A sample of chlorine has a volume of 1.26 L when the temperature is 107.4 degC. Assuming that the pressure is constant, at what temperature will the volume be 1058 mL? (Express your answer in Kelvins.)

OpenStudy (anonymous):

can some one help with my question

OpenStudy (abmon98):

As you increase the temperature molecules move more rapidly and would expand increasing volume. They have a directly proportional relation. if temperature increase volume increases and if temperature decreases volume decreases Use T1/T2=V1/V2 We can change Degree celsius by adding 273 and change 1.26 L to ml by multiplying by 1000 (107.4+273)/T2=1260/1058